Solve by the substitution method.
4x+7y = 35
-3x+y = 30
Can someone please help me with this.

Answer by oscargut(2103)

4x+7y = 35
-3x+y = 30
y=3x+30 (from 2nd eq)
then
using it at 1st eq we have
4x+7(3x+30)=35
25x+210=35
25x=-175
x=-7
y=3(-7)+30=9
Answer: x=-7,y=9